# Value Gradients with Action Adaptive Search Trees in Continuous (PO)MDPs

### v1

@@ B.1 AGMCTS Implementation Details
B.2 D-Continuous Light Dark
@@
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $x_{g}$, but the computed action has an added noise of $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$. In our experiments we used $\sigma_{r}=0.1$ for parameter tuning, and additionally evaluated for $\sigma_{r}=0.2,0.3$.
The implementation of D-Continuous Light Dark was written using POMDPs.jl [?]. The gradient function $\nabla_{a}r$ was equal 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [?]. Enzyme.jl provided extremely efficient gradients, achieving the execution times almost as close to the original functions. We did not care to optimize for making primal and dual calls, which might have allowed to save some computation time. However, it seems that it would have had a relatively minor effect as the heaviest computation was calculating the belief transition importance ratio, which due to it being a ratio of two products, cannot be MC approximated without introducing bias. Therefore, we iterated over all belief particles for its computation. Yet, in more complex problem where the dynamics are more complex, it might prove beneficial to compute gradients in the same pass as the transition likelihood wherever it can be done.
B.3 Evaluations
Parameter Notation Explanation
$c$ UCT exploration bonus
$k_{a}$ Action progressive widening (factor)
$\alpha_{a}$ Action progressive widening (power)
$k_{o}$ Observation progressive widening (factor)
$\alpha_{o}$ Observation progressive widening (power)
$\eta_{Adam}$ Adam step size (i.e. learning rate)
$T_{d_{a}}$ Action update distance threshold
Table 2: Hyperparameters optimized by CE in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions.
@@ B.3 Evaluations
Table 3: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light Dark domain, for $D$ dimensions, rounded to $3$ significant digits after the decimal point.
We evaluated the algorithms on the D-Continuous Light Dark domain for $D=2,3,4$ dimensions. PFT-DPW and AGMCTS were given a budget of 500 simulations. The particle count during planning was set to be dependent on the dimension:
$J_{D=2}=256,\quad J_{D=3}=512,$
$J_{D=4}=1024.$ (108)
@@
$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,$
$J^{\text{PF}}_{D=4}=8192.$ (109)

### v3

@@ B.3 Gradient Optimization
Since action updates are more costly than computing the gradient estimates, especially when not performing a linearized update of the importance weights, we set a threshold action distance for action update in some domains. After each gradient iteration, we update an accumulator of the gradient steps $a_{t}^{acc}$. Only after passing a threshold distance $\lVert{a^{acc}-a}\rVert>T_{d_{a}}^{\min}$, we update $a\leftarrow a^{acc}$. This limits the number of action updates, and acts like a convergence criterion. To ensure controllable step-sizes, we also set a max-norm threshold on the final action update $\lVert{a^{\prime}-a}\rVert<T_{d_{a}}^{\max}$.
For the Adam optimizer, we used the standard hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$. The learning rate $\eta_{Adam}$, alongside $T_{d_{a}}^{\min}$ in domains it was not set to 00, were determined by cross entropy optimization (CE) described later, in conjunction with the optimization over the other algorithm hyperparameters.
B.4 Thresholds For Adding/Deleting State Nodes
To save computational overhead, we would like to ensure that we have relevant samples after several action updates. As a relevance heuristic, we delete child state $s^{\prime}$ from $C(s,a)$ if $\omega_{a}(s^{\prime})<T_{\omega}^{del}$ after action update. Additionally, if $\omega_{a}(s^{\prime})<T_{\omega}^{add}$ for all $s^{\prime}\in C(s,a)$, we force to sample a new posterior node, overriding the progressive widening limitation. This ensures that we do not waste computational resources on states with low contribution to the estimator $\hat{Q}(s,a)$.
Due to the possibility of deleting posterior children, care must be taken when updating visitation counters since they do not monotonically increment anymore. For this reason, equations (13), (14), (11), (15), (16) are written using previous and updated visitation counts.
Appendix C Experiments
C.1 Experimental Method
Table 1: Hyperparameters optimized by CE in the evaluations of each algorithm in the D-Continuous Light-Dark, Mountain Car, Hill Car and Lunar Lander domains.
Parameter Notation Explanation Algorithms Tuning
$c$ UCT exploration bonus All CE
$k_{a}$ Action prog. widening factor All CE
$\alpha_{a}$ Action prog. widening power All CE
$k_{o}$ Obs. prog. widening factor All CE
$\alpha_{o}$ Obs. prog. widening power All CE
$\eta_{Adam}$ Adam step size (learning rate) AGMCTS CE
$T_{d_{a}}^{\min}$ Dist. threshold AGMCTS Manual/CE
$J$ No. particles AGMCTS/(PFT-)DPW Manual
$K_{rollout}$ No. states for rollout AGMCTS/(PFT-)DPW Manual
$K_{opt}$ No. grad. iterations AGMCTS Manual
$T_{d_{a}}^{\max}$ Max. action update step AGMCTS Manual
$T_{\omega}^{add}$ Imp. ratio add threshold AGMCTS Manual
$T_{\omega}^{del}$ Imp. ratio delete threshold AGMCTS Manual
$K_{\text{child}}^{\min}$ Action update min. children AGMCTS Manual
$K_{\text{child}}^{visits}$ Action update every mod $\text{mod }k$ AGMCTS Manual
$K_{b}^{\nabla}$ No. states for ^$\hat{\nabla}_{a}\log p_{T}$ AGMCTS Manual
$K_{O}^{\nabla}$ No. obs. branches for ^$\hat{\nabla}_{a}\hat{Q}$ AGMCTS Manual
$n^{sims}$ Simulation budget All Manual
The different hyperparameters and their explanations defined by each algorithm are summarized in Table 1.
In high-level, the experiments were conducted for each domain in the following manner:
  1. 1.
    Constant hyperparameters were defined for each algorithm, including simulation budget $n^{sims}$, particle count $J$, and rollout policy.
  2. 2.
    A cross-entropy (CE) optimization was performed over the UCB, DPW and step size (where applicable) for each algorithm.
  3. 3.
    For the best performing hyperparameters of each algorithm, we ran 1000 simulations of the algorithm with the same seeds between algorithms, at the multiples $[10^{-1},10^{-0.75},10^{-0.5},10^{-0.25},1.0]$ of the simulation budget $n^{sims}$.
The simulation jobs were run on a distributed cluster of 8 machines with different cores and configurations, using the Julia programming language with Distributed.jl multi-processing tools. Each simulation was run on a single core, without any parallelization in the solver itself. The processors used were:
  1. 1.
    Ryzen 7 3700X 8-Core (16 threads) Processor.
  2. 2.
    Intel i7-6820HQ 4-Core (8 threads) Processor.
  3. 3.
    Intel i7-8550U 4-Core (8 threads) Processor.
  4. 4.
    Intel E3-1241 v3 4-Core (8 threads) Processor.
  5. 5.
    Intel E5-1620 v4 4-Core (8 threads) Processor.
  6. 6.
    Intel i9-7920X 12-Core (24 threads) Processor.
  7. 7.
    Intel i9-11900K 8-Core (16 threads) Processor.
  8. 8.
    Intel i9-12900K 12-Core (24 threads) Processor.
C.1.1 Simulation Budget
Since POMCPOW is a state-trajectory algorithm, in the POMDP scenarios we measured empirically the runtime of PFT-DPW at $n^{sims}$ and $J$, and set a larger simulation budget for POMCPOW $n^{sims,\text{POMCPOW}}$ to match the runtime of PFT-DPW.
C.1.2 Rollout Policy
In order to save computation time and not compute a partially observable rollout, the rollout computation for a new particle belief was done by drawing $K_{rollout}$ particles from the belief, and computing the rollout policy based on the mean state, while calculating the rollout estimate based on the mean of the returns of the $K_{rollout}$ particles. For POMCPOW, since every new node is expanded initially only with a single particle, the rollout was computed with that particle. This is a disadvantage for POMCPOW as it made the rollout returns less meaningful, however it seemed to have achieved consistently better results than PFT-DPW despite that.
C.1.3 Cross-Entropy (CE) Optimization
The algorithm hyperparameters were first determined by cross entropy optimization (CE) [40], largely following [41], using Gaussian distributions over the continuous DPW hyperparameters:
  1. 1.
    UCT exploration bonus $c$.
  2. 2.
    Action progressive widening parameters $k_{a}$ and $\alpha_{a}$.
  3. 3.
    Observation progressive widening parameters $k_{o}$ and $\alpha_{o}$.
For AGMCTS, we also optimized over the Adam learning rate $\eta_{Adam}$, and the action update distance threshold $T_{d_{a}}^{\min}$ only in the Light-Dark domains.
We used 150 parameter samples, 40 simulations for each parameter sample to determine its mean return, used 30 elite samples to fit the new distribution, with smoothing the new distribution’s parameters:
where we chose $\alpha_{\mu}=0.8,\alpha_{\Sigma}=0.5$. When fine-tuning parameters based on previous runs, we used the previous mean and covariance as the starting point for the new CE optimization, with 80 parameter samples and 20 elite samples.
We performed 50 CE iterations for each algorithm in each dimension, with reasonable initial guesses for the starting parameters based on the literature and manual experiments. For some domains, we tuned the parameters of AGMCTS based on the best performing parameters of PFT-DPW, and kept the shared parameters constant throughtout the CE iterations of AGMCTS to limit it search space. The hyperparameters chosen were those from the CE iteration achieving the maximal mean of mean returns over the 30 elite samples. After choosing the parameters, we reported the mean return over 1000 scenarios, with the standard error of the mean.
C.2 D-Continuous Light-Dark
(a) Middle of scenario, $t=5$.
(b) End of scenario, $t=6$.
Figure 4: AGMCTS at 2-Continuous Light-Dark. The agent’s current state is the red dot, the current belief particles in orange, the next state in black, and the next belief particles in gray, the next observation is the green dot. The goal is the blue ring centered at $(0,2.5)$, $b_{0}$ is the red ring centered at $(0,0)$, and the beacon is the green star at $(2.5,0)$. The reward function for the posterior state is the blue-red heatmap drawn in the background. The shaded area is where the observation noise is $\sigma_{O}>5$. The chosen action branch of AGMCTS is drawn as the blue arrows, getting darker in hue with each action update. We can see that the optimized actions point closer to the goal from the current belief’s mean.
C.2.1 Problem Description
@@
To simulate more difficult domains, we include a noisy rollout policy. The rollout heads directly to $x_{g}$, but the computed action has an added noise of $\mathcal{N}(\mathbf{0},\sigma_{r}^{2}\mathbf{I})$. In our experiments we set $\sigma_{r}=0.1$.
The implementation of D-Continuous Light-Dark was written using POMDPs.jl [32]. The gradient function $\nabla_{a}r$ was equal to 00 in our setting, and $\nabla_{a}\log p_{T}$ was calculated using Enzyme.jl [33].
In this domain, it is straightforward to compute the transition density:
$p_{T}(x^{\prime}\mid x,a)=\frac{1}{(2\pi\sigma_{T}^{2})^{D/2}}\exp\left(-\frac{\lVert{x^{\prime}-(x+a)}\rVert^{2}}{2\sigma_{T}^{2}}\right).$ (156)
Enzyme.jl provided extremely efficient gradients, achieving the execution times almost as close to the original functions. However, it seems that it would have had a relatively minor effect as the heaviest computation was calculating the belief transition importance ratio, which due to it being a ratio of two products, cannot be MC approximated without introducing bias. Therefore, we iterated over all belief particles for its computation. Yet, in more complex problem where the dynamics are more complex, it might prove beneficial to compute gradients in the same pass as the transition likelihood wherever it can be done.
C.2.2 Evaluation
Table 2: Hyperparameters used in the evaluations of each algorithm for the D-Continuous Light-Dark domain, for $D=2,3,4$ dimensions, rounded to $3$ significant digits after the decimal point.
@@ C.2.2 Evaluation
We evaluated the algorithms on the D-Continuous Light-Dark domain for $D=2,3,4$ dimensions. PFT-DPW and AGMCTS were given a budget of $n^{sims}=500$ simulations. The particle count during planning was set to be dependent on the dimension:
$J_{D=2}=256,\quad J_{D=3}=512,\quad J_{D=4}=1024.$ (157)
@@
$J^{\text{PF}}_{D=2}=2048,\quad J^{\text{PF}}_{D=3}=4096,\quad J^{\text{PF}}_{D=4}=8192.$ (158)